OpenStudy (anonymous):

What is the area of this square in square centimeters? A. 25 square centimeters B. 2500 square centimeters C. 25,000 square centimeters D. 250,000 square centime

OpenStudy (anonymous):

@MD152727

OpenStudy (shamim):

1mm=?cm

OpenStudy (shamim):

50mm=?cm

OpenStudy (shamim):

U know area of a square=side^2

OpenStudy (shamim):

U know 1cm=10mm Right?

OpenStudy (shamim):

So 1mm=?cm

OpenStudy (shamim):

Response plz !

OpenStudy (shamim):

U should not expect answer only

OpenStudy (taric25):

\[50*10^-3 * 50*10^-3=0.0025\] Now you have the answer in square meters. Convert so square centimeters. \[0.0025*(\frac{ 100 }{ 1 })^2 = 25 \] The answer is A, 25 square centimeters.

OpenStudy (shamim):

U shpuld not tell the tesult directly

OpenStudy (shamim):

Should